# Variations on five-dimensional sphere packings

### v1

@@ Abstract.
We analyze Szöllősi’s recent construction of a five-dimensional kissing configuration and produce a new such configuration, the fourth to be discovered. We construct five-dimensional sphere packings from these configurations, which augment Conway and Sloane’s list of conjecturally optimal packings, and we also construct a new kissing configuration in nine dimensions. None of these constructions improves on the known records, but they provide geometrically distinct constructions achieving these records.
@@ 1. Introduction
In dimensions where no proof is known, it is natural to wonder how confidently we can identify the optimal packings. Conway and Sloane [4] took a major step in this direction by classifying how packings can be obtained by stacking layers, and they gave a conjectural list of optimal sphere packings in dimensions up to nine. (Above nine dimensions, this approach fails, which illustrates the increasing complexity of the sphere packing problem.) Their list does not contain every possible densest packing, because one can always make local perturbations without changing the global density, but they conjectured that it contains all the “tight” packings. Kuperberg [9] disproved their specific formulation of tightness and suggested a replacement in terms of weak recurrence (Conjecture 15 in [9]). With this amended formulation, Conway and Sloane’s list of packings seemed plausibly complete, as did the analogous classification of kissing configurations by Cohn, Jiao, Kumar, and Torquato [2]. However, in 2023 Szöllősi [19] found a new five-dimensional kissing configuration, which could be extended to a new five-dimensional sphere packing (see the acknowledgements in [19]). This example showed that even in a few as five dimensions, our understanding of optimal sphere packings was incomplete.
@@
In this paper, we put Szöllősi’s construction in a broader context. We construct another five-dimensional kissing configuration, which brings the list of known examples to four, and we extend these kissing configurations to a family of five-dimensional sphere packings. It is conceivable that our list completes Conway and Sloane’s list in five dimensions, but it is difficult to rule out further surprises. The two most noteworthy of the new sphere packings are uniform, meaning that their symmetry groups act transitively on the spheres. Only four such packings had previously been known, namely the $D_{5}$ root lattice and three packings constructed by Leech [10].
@@
We have not had any luck applying our techniques to six or seven dimensions. We expect that further kissing configurations and packings remain to be discovered in these dimensions, but doing so may require a new idea. Instead, we construct a new example in nine dimensions, which again does not set a new record but provides a geometrically distinct way to achieve the known record.
@@
All of these constructions can be understood through modifying layers in packings. We will describe kissing configurations and packings in terms of the locations of the sphere centers. In other words, a kissing configuration consists of an arrangement of points on the surface of a sphere such that each pair of distinct points forms an angle of at least $\pi/3$ with the center of the sphere, and a sphere packing consists of an arrangement of points in Euclidean space with a certain minimal distance (namely, twice the sphere radius). Our altered configurations will remove the points on a hyperplane and replace them with modified points. The simplest case is the three-dimensional kissing configuration from the face-centered cubic lattice, in which it is well known that one can replace a triangle of points as shown in Figure 1.1.
@@
In the remainder of this paper, we analyze the five-dimensional kissing configurations and construct our new configuration (Section 2), examine what goes wrong in six dimensions (Section 3), extend these kissing configurations to sphere packings (Section 4), and conclude with our construction in nine dimensions (Section 5).
@@ 2. Five-dimensional kissing configurations
$v\mapsto v-2\langle v,w\rangle w,$
@@
Table 2.1 lists the inner products that occur in these configurations and their multiplicities, while Table 2.2 list coordinates for the points themselves. The symmetry groups of $D_{5}$, $L_{5}$, $Q_{5}$, and $R_{5}$ have orders $3840$, $384$, $240$, and $48$, respectively. Specifically, the symmetries of $D_{5}$ are generated by signed permutations of the coordinates, and those of $L_{5}$ are signed permutations that fix the fifth coordinate. The symmetries of $Q_{5}$ are generated by the $120$ symmetries of the $A_{4}$ cross section (which are isomorphic to the permutation group $S_{5}$) and reflection across the hyperplane, and the same is true for $R_{5}$ except that there are only $24$ symmetries for its modified cross section (isomorphic to the permutation group $S_{4}$).
@@ 4. From kissing configurations to sphere packings
The local analysis from Figure 4.1 leads to a complete classification of edge-to-edge tilings with the tiles $\Delta_{1}$ and $\Delta_{2}$. Among these possibilities, those shown in Figure 4.2 are the only ones that have kissing configuration $Q_{5}$ or $R_{5}$ everywhere. However, other combinations of kissing configurations are possible, such as tilings that have layers of types $a$ and $b$, in which case some spheres have kissing configuration $Q_{5}$ or $R_{5}$ and some have kissing configuration $D_{5}$ or $L_{5}$, and one can combine all four kissing configurations by varying the coloring condition in Table 4.1 among the different spheres in a packing.
Conjecture 4.3.
Of course Szöllősi’s discovery is cause for skepticism about any such classification. We have no great confidence in Conjecture 4.3, but it describes the current state of our knowledge. Even packings that do not obviously fit into this framework can often be reformulated in this way. For example, the periodic packing described in the acknowledgements of [19] is equivalent to the uniform packing with kissing configuration $Q_{5}$ from Figure 4.2.
@@
This four-coloring construction of five-dimensional packings works equally well to produce six-dimensional sphere packings from valid four-colored point configurations in $\mathbb{R}^{3}$. One can construct such configurations from face-to-face tilings of $\mathbb{R}^{3}$ with irregular tetrahedra and octahedra that have $\Delta_{1}$ faces, as shown in Figure 4.3. These polyhedra have many lovely properties. For example, one octahedron and four tetrahedra can be assembled to form a larger tetrahedron (with doubled edge lengths, and with the tetrahedra nestled in the corners of the larger tetrahedron), while six octahedra and eight tetrahedra can form a larger octahedron (again with doubled edge lengths and the octahedra nestled in the corners). Iterating this procedure yields a tiling of all of space, which can be colored to produce the $E_{6}$ packing. Furthermore, other tilings yield all the six-dimensional sphere packings from [4]. However, we do not obtain any new six-dimensional sphere packings. Somehow the five-dimensional case has additional flexibility, which we have been unable to replicate in six or seven dimensions.

### v2

@@ Abstract.
We analyze Szöllősi’s recent construction of a conjecturally optimal five-dimensional kissing configuration and produce a new such configuration, the fourth to be discovered. We construct five-dimensional sphere packings from these configurations, which augment Conway and Sloane’s list of conjecturally optimal packings. We also construct a new kissing configuration in nine dimensions. None of these constructions improves on the known records, but they provide geometrically distinct constructions achieving these records.
@@ 1. Introduction
In dimensions where no proof is known, it is natural to wonder how confidently we can identify the optimal packings. Conway and Sloane [5] took a major step in this direction by classifying how packings can be obtained by stacking layers, and they gave a conjectural list of optimal sphere packings in dimensions up to nine. (Above nine dimensions, this approach fails, which illustrates the increasing complexity of the sphere packing problem.) Their list does not contain every possible densest packing, because one can always make local perturbations without changing the global density, but they conjectured that it contains all the “tight” packings. Kuperberg [10] disproved their specific formulation of tightness and suggested a replacement in terms of weak recurrence (Conjecture 15 in [10]). With this amended formulation, Conway and Sloane’s list of packings seemed plausibly complete, as did the analogous classification of kissing configurations by Cohn, Jiao, Kumar, and Torquato [3]. However, in 2023 Szöllősi [20] found a new five-dimensional kissing configuration, which could be extended to a new five-dimensional sphere packing (see the acknowledgements in [20]). This example showed that even in as few as five dimensions, our understanding of optimal sphere packings was incomplete.
@@
In this paper, we put Szöllősi’s construction in a broader context. We construct another five-dimensional kissing configuration, which brings the list of known examples to four, and we extend these kissing configurations to a family of five-dimensional sphere packings. It is conceivable that our list completes Conway and Sloane’s list in five dimensions, but it is difficult to rule out further surprises. The two most noteworthy sphere packings in the new family are uniform, which means their symmetry groups act transitively on the spheres. Only four such packings had previously been analyzed, namely the $D_{5}$ root lattice and three packings constructed by Leech [11]. We prove that these six uniform packings are the only uniform packings that contain the known kissing configurations locally.
The unique uniform packing with Szöllősi’s kissing configuration was previously discovered by Andreanov and Kallus [1] as $J_{5,5}$ in their classification of dense $2$-periodic packings (i.e., unions of two translates of a lattice). In fact, their discovery predates Szöllősi’s, although they did not recognize that the kissing configuration was novel. The remaining uniform packing, with our new kissing configuration, is $4$-periodic. Viewing these uniform packings as $m$-periodic with $m$ small makes it easier to analyze their symmetry groups.
@@
We have had no luck applying our techniques to six or seven dimensions. We expect that further kissing configurations and packings remain to be discovered in these dimensions, but doing so may require a new idea. Instead, we construct a new example in nine dimensions, which again does not set a new record but provides a geometrically distinct way to achieve the known record.
@@
All of these constructions can be understood through modifying layers in packings. We will describe kissing configurations and packings in terms of the locations of the sphere centers. In other words, a kissing configuration consists of an arrangement of points on the surface of a sphere such that each pair of distinct points forms an angle of at least $\pi/3$ with the center of the sphere, and a sphere packing consists of an arrangement of points in Euclidean space with a certain minimal distance (namely, twice the sphere radius). Our altered configurations will remove the points on a hyperplane and replace them with modified points. The simplest case is the three-dimensional kissing configuration of the face-centered cubic lattice, in which it is well known that one can replace a triangle of points as shown in Figure 1.1.
@@
In the remainder of this paper, we analyze the five-dimensional kissing configurations and construct our new configuration (Section 2), examine what goes wrong in six dimensions (Section 3), extend these kissing configurations to sphere packings (Section 4), and conclude with our construction in nine dimensions (Section 5). We also provide computer code for verifying certain assertions through DSpace@MIT at https://hdl.handle.net/1721.1/157699.
@@ 2. Five-dimensional kissing configurations
(2.1) $v\mapsto v-2\langle v,w\rangle w,$
@@
Table 2.1 lists the inner products that occur in these configurations and their multiplicities, while Table 2.2 lists coordinates for the points themselves. The symmetry groups of $D_{5}$, $L_{5}$, $Q_{5}$, and $R_{5}$ have orders $3840$, $384$, $240$, and $48$, respectively. Specifically, the symmetries of $D_{5}$ are the signed permutations of the coordinates, and those of $L_{5}$ are the signed permutations that fix the fifth coordinate. The symmetries of $Q_{5}$ are generated by the permutations of the coordinates and reflection across the hyperplane, and the symmetries of $R_{5}$ are generated by the permutations of the coordinates that fix the fifth coordinate and reflection across the hyperplane.
@@ 4.1. Packings that fiber over $D_{3}$
The local analysis from Figure 4.1 leads to a complete classification of edge-to-edge tilings with the tiles $\Delta_{1}$ and $\Delta_{2}$. Among these possibilities, those shown in Figure 4.2 are the only ones that have kissing configuration $Q_{5}$ everywhere or $R_{5}$ everywhere. However, other combinations of kissing configurations are possible, such as tilings that have layers of types $a$ and $b$, in which case some spheres have kissing configuration $Q_{5}$ or $R_{5}$ and some have kissing configuration $D_{5}$ or $L_{5}$. One can combine all four kissing configurations in a single packing by varying the coloring condition in Table 4.1 among the different spheres in a packing.
Conjecture 4.2.
Of course Szöllősi’s discovery is cause for skepticism about any such classification. We have no great confidence in Conjecture 4.2, but it describes the current state of our knowledge.
4.2. Uniform packings
@@
Table 4.2. The two uniform packings from Figure 4.2, with the left one expressed as a 2-periodic packing above and the right one as a 4-periodic packing below.
Lattice basis Translation vectors
$Q_{5}$ uniform packing
$v_{0}$ ${}=(1,-1,0,0,0)$ $x_{0}$ ${}=(0,0,0,0,0)$
$v_{1}$ ${}=(0,1,-1,0,0)$ $x_{1}$ ${}=(0,0,0,-1,-1)$
$v_{2}$ ${}=(0,0,1,-1,0)$
$v_{3}$ ${}=(0,0,0,1,-1)$
$v_{4}$ ${}=(0.8,0.8,0.8,0.8,0.8)$
$R_{5}$ uniform packing
$v_{0}$ ${}=(1,-1,0,0,0)$ $x_{0}$ ${}=(0,0,0,0,0)$
$v_{1}$ ${}=(0,1,-1,0,0)$ $x_{1}$ ${}=(0,0,0,1,-1)$
$v_{2}$ ${}=(0,0,1,-1,0)$ $x_{2}$ ${}=(0,0,-1,-1,0)$
$v_{3}$ ${}=(-0.5,-0.5,-0.5,-0.5,2)$ $x_{3}$ ${}=(0,0,0,-1,-1)$
$v_{4}$ ${}=(0.8,0.8,0.8,0.8,0.8)$
Given the set $\mathcal{P}$ of sphere centers in a periodic sphere packing in $\mathbb{R}^{n}$, let
$\Lambda=\{y\in\mathbb{R}^{n}:\mathcal{P}=\mathcal{P}+y\}$
be its translational symmetry lattice. Then $\mathcal{P}$ is the union of finitely many translates of $\Lambda$. It is called $m$-periodic if it consists of $m$ translates of $\Lambda$, and our choice of $\Lambda$ ensures that it cannot be written as the union of fewer translates of any other lattice (i.e., $m$ is minimal).
The uniform packing from Figure 4.2 with kissing configuration $Q_{5}$ is $2$-periodic, and the one with kissing configuration $R_{5}$ is $4$-periodic, with lattice bases and translation vectors shown in Table 4.2. Andreanov and Kallus [1] showed that no $2$-periodic sphere packing in $\mathbb{R}^{5}$ can be denser than the $D_{5}$ root lattice, and they obtained a complete list of the densest such packings. In the process they discovered the $Q_{5}$ uniform sphere packing, which also occurs in the acknowledgements of [20]. No such optimality theorem or classification is known beyond $2$-periodic packings. As for the remaining uniform packings, the $D_{5}$ root lattice is of course $1$-periodic, and the three $L_{5}$ uniform packings are $2$-periodic, $3$-periodic, and $4$-periodic (see [5]).
4.3. Symmetries of uniform packings
We will show in this subsection that each symmetry of the $Q_{5}$ or $R_{5}$ kissing configuration extends to a symmetry of the corresponding uniform packing that fixes a point. The symmetry group of the kissing configuration is the stabilizer of a point, and the semidirect product of this group and the translational symmetry lattice has index $m$ in the symmetry group of the uniform packing when the packing is $m$-periodic (i.e., $m=2$ for $Q_{5}$ and $m=4$ for $R_{5}$).
We enumerate the symmetries as follows. Each symmetry of a subset $\mathcal{P}$ of $\mathbb{R}^{n}$ is an affine linear map $x\mapsto Ax+b$ with $A$ orthogonal. If $\Lambda=\{y\in\mathbb{R}^{n}:\mathcal{P}=\mathcal{P}+y\}$ as in the previous subsection, then $A$ must preserve $\Lambda$. To see why, note that if $x\mapsto Ax+b$ is a symmetry of $\mathcal{P}$ and $y\in\Lambda$, then
$\{Ax+b:x\in\mathcal{P}\}=\mathcal{P}=\mathcal{P}+y=\{A(x+y)+b:x\in\mathcal{P}\}=\{Ax+b+Ay:x\in\mathcal{P}\},$
from which it follows that $Ay\in\Lambda$, as desired. Thus, determining the symmetries of $\mathcal{P}$ amounts to determining the symmetries of $\Lambda$ and which translation vectors $b$ are compatible with them.
For the $Q_{5}$ uniform packing, the underlying lattice $\Lambda$ is the orthogonal direct sum of the $A_{4}$ root lattice and a one-dimensional lattice (see Table 4.2 for a lattice basis). The symmetry group $G$ of $\Lambda$ is isomorphic to $S_{5}\times C_{2}^{2}$, where the symmetric group $S_{5}$ permutes the five coordinates and the cyclic factors $C_{2}$ are generated by scalar multiplication by $-1$ and the reflection $s$ across the hyperplane perpendicular to the fifth lattice basis vector $v_{4}$ shown in Table 4.2. The reflection $s$ is the same reflection as in equation (2.1), and it corresponds to a reflection across a horizontal line in Figure 4.2. Let $H$ be the subgroup of $G$ generated by $S_{5}$ and $s$ (the symmetry group of the kissing configuration $Q_{5})$, let $H_{0}=H$, and let $H_{1}=(-1)H$ be the coset obtained through scalar multiplication by $-1$. Then a case analysis shows that the symmetry group of the $Q_{5}$ uniform packing is
$\{x\mapsto Ax+x_{i}+y:A\in H_{i},\ i\in\{0,1\},\ y\in\Lambda\},$
where the translation vectors $x_{0}$ and $x_{1}$ are specified in Table 4.2. Each symmetry of $\Lambda$ corresponds to a unique translation vector, modulo $\Lambda$.
For the $R_{5}$ uniform packing, the underlying lattice $\Lambda$ is the orthogonal direct sum of the $A_{3}$ root lattice and two differently scaled one-dimensional lattices (again see Table 4.2 for a lattice basis). The symmetry group $G$ of $\Lambda$ is isomorphic to $S_{4}\times C_{2}^{3}$, where the symmetric group $S_{4}$ permutes the first four coordinates and the cyclic factors $C_{2}$ are generated by scalar multiplication by $-1$ and the reflections $r$ and $s$ across the hyperplanes perpendicular to the fourth and fifth lattice basis vectors $v_{3}$ and $v_{4}$ shown in Table 4.2. The reflections $r$ and $s$ correspond to reflections across vertical and horizontal lines in Figure 4.2, respectively. Let $H$ be the subgroup of $G$ generated by $S_{4}$ and $s$ (the symmetry group of the kissing configuration $R_{5})$, and let $H_{0}=H$, $H_{1}=(-r)H$, $H_{2}=rH$, and $H_{3}=(-1)H$ be its cosets, where as above $(-1)$ denotes scalar multiplication by $-1$ and $-r$ denotes the composition of $(-1)$ and $r$. Then a case analysis shows that the symmetry group of the $R_{5}$ uniform packing is
$\{x\mapsto Ax+x_{i}+y:A\in H_{i},\ i\in\{0,1,2,3\},\ y\in\Lambda\},$
where again the translation vectors $x_{i}$ are specified in Table 4.2. As in the case of $Q_{5}$, each symmetry of $\Lambda$ corresponds to a unique translation vector, modulo $\Lambda$.
4.4. Higher dimensions
@@
The four-coloring construction of five-dimensional packings from Section 4.1 works equally well to produce six-dimensional sphere packings from valid four-colored point configurations in $\mathbb{R}^{3}$. One can construct such configurations from face-to-face tilings of $\mathbb{R}^{3}$ with irregular tetrahedra and octahedra that have $\Delta_{1}$ faces, as shown in Figure 4.3. These polyhedra have many lovely properties. For example, one octahedron and four tetrahedra can be assembled to form a larger tetrahedron (with doubled edge lengths, and with the tetrahedra nestled in the corners of the larger tetrahedron), while six octahedra and eight tetrahedra can form a larger octahedron (again with doubled edge lengths and the octahedra nestled in the corners). Iterating this procedure yields a tiling of all of space, which can be colored to produce the $E_{6}$ packing. Furthermore, other tilings yield all the six-dimensional sphere packings from Section 6 of [5]. However, we do not obtain any new six-dimensional sphere packings. Somehow the five-dimensional case has additional flexibility, which we have been unable to replicate in six or seven dimensions.
@@ 5. Nine-dimensional kissing configurations
Acknowledgements
We are grateful to Yoav Kallus for pointing out that the kissing configuration identified by Szöllősi in [20] had previously occurred as the kissing configuration of a packing discovered in [1].